OpenStudy (anonymous):

find the point of intersection of y=x^2 and y=12-2x^2

OpenStudy (saifoo.khan):

Equate them both.

OpenStudy (saifoo.khan):

y=x^2 and y=12-2x^2 x^2 = 12 - 2x^2 Solve.

OpenStudy (anonymous):

x^2 = 12 - 2x^2 x^2 +2x^2 = 12 continue it

OpenStudy (anonymous):

@ yasmin95

OpenStudy (anonymous):

@yasmin95, are you ok now at that point? can you follow now?

OpenStudy (anonymous):

once you add the coefficient of x^2 then you can solve for x and y that will be the point of intersection

OpenStudy (anonymous):

__?x^2=12 or x^2 +2x^2 = 12 (1+2)x^2=12

OpenStudy (anonymous):

once you find x, sub this x to any of this y=x^2 and y=12-2x^2